OpenStudy (anonymous):

Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

mathslover (mathslover):

is their any equation given?

OpenStudy (saifoo.khan):

Hey @SchoolFreak , we meet again. ;)

OpenStudy (anonymous):

MathsLover this is for the last one we were doing

mathslover (mathslover):

ok you mean the equation is : -3x+2y = 8

OpenStudy (anonymous):

@saifoo.khan lol yup

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok so first of all the answer was that in question itself see first of all : find the y intercept : put x = 0

mathslover (mathslover):

You got that earlier y = 4 . right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

-3(0)+2y =8 2y = 8 y =4

mathslover (mathslover):

hence when x = 0 then y =4

mathslover (mathslover):

now ... put x = 2 ,, y = 7 , right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

-3(2)+2y = 8 -6+2y = 8 2y=14 y=7 (as described earlier)

mathslover (mathslover):

now .. we have some coordinates , (0,4) and (2,7)

mathslover (mathslover):

mark them on graph .. first

OpenStudy (anonymous):

k

OpenStudy (anonymous):

y = mx + b. <=== arent we suppose to use this for this question

mathslover (mathslover):

no ... @SchoolFreak since we have to draw the graph why to use that? just mark the coordinates on the graph and join them ..

mathslover (mathslover):